OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57

OpenStudy (texaschic101):

7x - 2y = 21 4x + y = 57 -->(2) -------------- 7x - 2y = 21 8x + 2y = 114 (result of multiplying by 2) -------------add 15x = 135 x = 135/15 x = 9